# Structured coalescents, coagulation equations and multi-type branching processes

### v2

@@ 2.3. Convergence
In this section we state the two main results concerning the convergence of the empirical measures $(\mu_{i}^{K})_{i\in[d]}$ as $K\to\infty$. Let us anticipate that our results will differ substantially depending on the parameter regime under consideration.
We start by introducing some notation. If $\eta_{1},\eta_{2}\in{\mathcal{M}}_{f}({\mathbb{R}}^{d})$ we denote by $\eta_{1}\star\eta_{2}$ their convolution, i.e.
@@
In addition, if $\eta\in{\mathcal{M}}_{f}({\mathbb{R}}^{d})$ and $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is integrable with respect to $\eta$, we set
@@
If the involved measures have discrete support, the integrals can, as usual, be understood as sums.
Now we state our main convergence result in the critical regime.
@@
Assume that $\gamma_{K}\to{c}$ as $K\to\infty$. If Assumption 2.3 holds, then $(\mu_{i}^{K})_{i\in[d]}$ weakly converges, as $K\to\infty$, to the solution of the $d$-dimensional discrete coagulation equation
$\partial_{t}u_{i}(t,{\boldsymbol{n}})$ $=\alpha_{i}\left(\frac{1}{2}u_{i}\star u_{i}(t,{\boldsymbol{n}})-\left<u_{i}(t),1\right>u_{i}(t,{\boldsymbol{n}})\right)+\sum_{j\in[d]\setminus\{i\}}w_{j,i}u_{j}(t,{\boldsymbol{n}})-w_{i,j}u_{i}(t,{\boldsymbol{n}}),$ (2.3)
@@
The dynamics of the coagulation equation is very intuitive. When two blocks coalesce, the resulting block configuration is the sum of the block configurations of the coalescing blocks; hence the convolution. Moreover, whenever a block with a given configuration merges with another block, we lose that configuration. The remaining sum is due to migration: we gain a block configuration in colony $i$ whenever a block with that configuration migrates from another colony to colony $i$; we lose a block configuration whenever a block with that configuration migrates from colony $i$ to another colony $i$. The initial condition is due to the fact that at time $t=0$ there are approximately $\beta_{i}K$ singletons in colony $i$.
The next theorem covers the large sampling case.
@@
Assume that $\gamma_{K}\to\infty$ as $K\to\infty$. If Assumption 2.3 holds, then $(\mu_{i}^{K})_{i\in[d]}$ weakly converges, as $K\to\infty$, to the weak solution of the $d$-dimensional continuous coagulation equation
$\partial_{t}u_{i}=\alpha_{i}\left(\frac{1}{2}u_{i}\star u_{i}-u_{i}\left<u_{i},1\right>\right)\ +\ \sum_{j\in[d]\setminus\{i\}}w_{ji}u_{j}-w_{ij}u_{i},\quad i\in[d],\,t>0$ (2.5)
$\lim_{t\to 0}\int\left(1-e^{-\left<\boldsymbol{\lambda},{\boldsymbol{x}}\right>}\right)u_{i}(t,{\rm d}{\boldsymbol{x}})=\lambda_{i}\beta_{i},\quad\boldsymbol{\lambda}=(\lambda_{i})_{i\in[d]}>\boldsymbol{0},\,i\in[d].$ (2.6)
@@
By weak solution, we mean that for $f\in C_{b}(\mathbb{R}_{+}^{d})$
$\langle u_{i}(t),f\rangle=\ \langle u_{i}(0),f\rangle$ $+\frac{\alpha_{i}}{2}\int_{0}^{t}(\langle(u_{i}\star u_{i})(s),f\rangle-2\ \langle u_{i}(s),1\rangle\ \langle u_{i}(s),f\rangle)\mathrm{d}s$
$+\sum_{j\in[d]\setminus\{i\}}\int_{0}^{t}(w_{j,i}\ \langle u_{j}(s),f\rangle-w_{i,j}\text{ }\langle u_{i}(s),f\rangle)\mathrm{d}s.$
@@ 2.4. Stochastic representation
In this section we provide stochastic representations to the coagulation equations obtained in the previous section. We provide here the derivation only on the large sampling case; the critical sampling case will be treated analogously in Appendix A. Therefore, throughout this section, we assume that $\gamma_{K}\to\infty$ as $K\to\infty$.
Let ${{\boldsymbol{u}}}$ be the solution of the continuum coagulation equation (2.5) under (2.6). Define
@@
A direct computation shows that ${\boldsymbol{v}}$ solves the multi-dimensional ODE
@@
We identify ${\boldsymbol{v}}$ as the Laplace exponent of the ${\boldsymbol{\psi}}$-CSBP, or equivalently the $d$-dimensional Feller diffusion ${{\boldsymbol{Z}}}$ solution of the SDE (see e.g. [37, Chap. 4.2, Eq. (4.4)])
@@
where ${\boldsymbol{B}}$ is a standard $d$-dimensional Brownian motion. Thus, for every ${\boldsymbol{x}}=(x_{i})_{i\in d]}$, we have
$\mathbb{E}_{{\boldsymbol{x}}}(e^{-{\boldsymbol{\lambda}}\cdot{\boldsymbol{Z}}_{t}})\ =\ e^{-\left<{{\boldsymbol{x}}},{\boldsymbol{v}}(t,{\boldsymbol{\lambda}})\right>}.$ (2.8)
Let $Q_{i}$ be the entrance law of the process from state $i$, i.e.
$\ \ Q_{i}(t,A)\ \coloneqq\ \lim_{x\to 0+}\frac{\mathbb{P}_{x{\boldsymbol{e}}_{i}}(Z_{t}\in A)}{x},\quad t>0.$
Using standard arguments (see e.g. [30, Thm. 3.13]), one can deduce from (2.8) that for $t>0$
@@
The solution of the coagulation equation (2.5) with initial condition (2.6) admits the stochastic representation
@@
Then, the solution of the coagulation equation (2.3) with initial condition (2.4) admits the stochastic representation
@@
In particular, the previous result holds for sufficiently large values of $c$. This aligns with the large sampling case, in which $c$ can be understood as being infinite.
The uniqueness of the solution to the coagulation equation (2.3) with initial condition (2.4) will be addressed in the Appendix; the assumption $d_{i}\geq 0$ will be not required.
2.5. Proof strategy
The remainder of the manuscript is devoted to the proofs of Theorems 2.4 and 2.5. In Section 3.1, we characterize the generator of the process of empirical measures and establish its convergence in a precise sense. In Section 3.4 we establish a coupling allowing us to compare the number of blocks in our multi-type coalescent to the number of blocks in Kingman’s coalescent; the resulting bounds will play an important role in later sections. In Section 4, we prove the tightness of the sequence of empirical measures and characterize its accumulation points. Finally, in Section 5, we complete the proofs of Theorems 2.4 and 2.5, offering further insight into the role of the initial condition in the large sampling regime (i.e. Eq. (2.4)).
3. Convergence of the generators and Kingman bounds
In this section we describe the action of the infinitesimal generator $A^{K}$ of the process of empirical measures ${\boldsymbol{\mu}}^{K}\coloneqq(\mu_{1}(t,\cdot),\ldots\mu_{d}(t,\cdot))_{t\geq 0}$ on a special class of functions. We will use $A^{K}$ to derive properties of the process ${\boldsymbol{\mu}}^{K}$. In the second half of this section we will prove the convergence as $K\to\infty$ of the generators $A^{K}$, in an appropriate sense.

### v3

@@ 2.4. Convergence
In this section, we state the two main results concerning the convergence of the empirical measures $(\mu_{i}^{K})_{i\in[d]}$ as $K\to\infty$. Let us emphasize from the outset that the qualitative behaviour of the limit depends strongly on the parameter regime under consideration.
We start by introducing some notation. If $\eta_{1},\eta_{2}\in\mathcal{M}_{f}(\mathbb{R}^{d})$, we denote by $\eta_{1}\star\eta_{2}$ their convolution, that is,
@@
Moreover, if $\eta\in\mathcal{M}_{f}(\mathbb{R}^{d})$ and $f:\mathbb{R}^{d}\to\mathbb{R}$ is integrable with respect to $\eta$, we set
@@
If the measures involved have discrete support, these integrals are, as usual, understood as sums.
We now extend the convergence result from the one-dimensional critical sampling regime to the multidimensional setting.
@@
Assume that $\gamma_{K}\to c$ as $K\to\infty$. If Assumption 2.3 holds, then $(\mu_{i}^{K})_{i\in[d]}$ converges weakly, as $K\to\infty$, to the solution of the $d$-dimensional discrete coagulation equation
$\partial_{t}u_{i}(t,\boldsymbol{n})$ $=\alpha_{i}\left(\frac{1}{2}u_{i}\star u_{i}(t,\boldsymbol{n})-\langle u_{i}(t),1\rangle\,u_{i}(t,\boldsymbol{n})\right)+\sum_{j\in[d]\setminus\{i\}}\big(w_{j,i}u_{j}(t,\boldsymbol{n})-w_{i,j}u_{i}(t,\boldsymbol{n})\big),$ (2.3)
@@
The dynamics of the coagulation equation are intuitive. When two blocks coalesce, the resulting block configuration is the sum of the configurations of the coalescing blocks, which explains the convolution term. Moreover, whenever a block with a given configuration merges with another block, that configuration is lost. The remaining terms account for migration: a block configuration is gained in colony $i$ when a block with that configuration migrates from another colony to $i$, and it is lost when a block migrates from colony $i$ to another colony. The initial condition reflects the fact that at time $t=0$ there are approximately $\beta_{i}K$ singletons in colony $i$.
The next theorem extends the convergence results to the large-sampling regime.
@@
Assume that $\gamma_{K}\to\infty$ as $K\to\infty$. If Assumption 2.3 holds, then $(\mu_{i}^{K})_{i\in[d]}$ converges weakly, as $K\to\infty$, to the weak solution of the $d$-dimensional continuous coagulation equation
$\partial_{t}u_{i}$ $=\alpha_{i}\left(\frac{1}{2}u_{i}\star u_{i}-u_{i}\langle u_{i},1\rangle\right)+\sum_{j\in[d]\setminus\{i\}}\big(w_{j,i}u_{j}-w_{i,j}u_{i}\big),\quad i\in[d],\;t>0,$ (2.5)
$\lim_{t\to 0}$ $\int_{\mathbb{R}_{+}^{d}}\left(1-e^{-\langle\boldsymbol{\lambda},\boldsymbol{x}\rangle}\right)u_{i}(t,\mathrm{d}\boldsymbol{x})=\lambda_{i}\beta_{i},\quad\boldsymbol{\lambda}=(\lambda_{i})_{i\in[d]}>\boldsymbol{0},\;i\in[d].$ (2.6)
@@
In the previous theorem, by a weak solution we mean that for $f\in C_{b}(\mathbb{R}_{+}^{d})$,
$\langle u_{i}(t),f\rangle$ $=\langle u_{i}(0),f\rangle+\frac{\alpha_{i}}{2}\int_{0}^{t}\big(\langle(u_{i}\star u_{i})(s),f\rangle-2\langle u_{i}(s),1\rangle\langle u_{i}(s),f\rangle\big)\,\mathrm{d}s$
$\quad+\sum_{j\in[d]\setminus\{i\}}\int_{0}^{t}\big(w_{j,i}\langle u_{j}(s),f\rangle-w_{i,j}\langle u_{i}(s),f\rangle\big)\,\mathrm{d}s.$
In the discrete setting underlying the critical sampling regime, weak convergence is defined analogously, but with the full class of bounded functions on $N]_{0}^{d}$ as test functions. Since this class contains indicator functions, the notions of weak and strong solutions coincide; for this reason, we do not distinguish between them in Theorem 2.4. Moreover, in both Theorem 2.4 and Theorem 2.5, existence of (weak) solutions is part of the conclusion.
The form of the initial condition in (2.6) is well known in the context of continuous-state branching processes (see Section 2.5). In the present setting, it reflects the fact that we start with $N_{K}$ blocks while rescaling total mass only by $K$, which leads to a singularity at $t=0$ as $K\to\infty$. In this regime, coalescence dominates the dynamics until the number of blocks reaches order $K$. We therefore expect migration to have little effect near time zero, and hence that in colony $i$ one has $\langle\boldsymbol{\lambda},\boldsymbol{x}\rangle\approx\lambda_{i}x_{i}$. This, in turn, suggests that as $t\to 0$,
$\int_{\mathbb{R}_{+}^{d}}\left(1-e^{-\langle\boldsymbol{\lambda},\boldsymbol{x}\rangle}\right)\mu_{i}^{K}(t,\mathrm{d}\boldsymbol{x})\;\approx\;\lambda_{i}\int_{\mathbb{R}_{+}^{d}}x_{i}\,\mu_{i}^{K}(t,\mathrm{d}\boldsymbol{x})\;\approx\;\lambda_{i}\frac{L_{i}^{K}(0)}{Ks_{K}}\;=\;\lambda_{i}\frac{L_{i}^{K}(0)}{N_{K}}.$
As $K\to\infty$, this expression converges to $\lambda_{i}\beta_{i}$, which is precisely the quantity appearing on the right-hand side of (2.6).
@@ 2.5. Stochastic representation
In this section, we generalize the stochastic representation established for the coagulation equation in the one-dimensional critical-sampling regime to the solutions of the coagulation equations obtained in the previous section. We derive these representations in the large-sampling regime; the critical-sampling regime is treated analogously in Appendix A. Throughout the remainder of this section, we therefore assume that $\gamma_{K}\to\infty$ as $K\to\infty$.
Assume that there exists a weak solution ${{\boldsymbol{u}}}=(u_{i})_{i\in[d]}$ of the continuum coagulation equation (2.5) under (2.6). Define
@@
A direct computation shows that ${\boldsymbol{v}}=(v_{i})_{i\in[d]}$ solves the multi-dimensional ODE
@@
We identify $\boldsymbol{v}$} as the Laplace exponent of the $\boldsymbol{\psi}$}-CSBP, or equivalently the $d$-dimensional Feller diffusion $\boldsymbol{Z}$}} solution of the SDE (see, for example, [45, Thm. 3])
@@
where $\boldsymbol{B}$} is a standard $d$-dimensional Brownian motion. Thus, for every ${\boldsymbol{x}}=(x_{i})_{i\in[d]}$, we have
$\mathbb{E}_{{\boldsymbol{x}}}[e^{-{\langle\boldsymbol{\lambda}},{\boldsymbol{Z}}(t)\rangle}]=e^{-\left<{{\boldsymbol{x}}},{\boldsymbol{v}}(t,{\boldsymbol{\lambda}})\right>}.$ (2.8)
Let $Q_{i}$ be the entrance law of the process starting from state $i$, i.e.
$Q_{i}(t,A)\coloneqq\lim_{x\to 0+}\frac{\mathbb{P}_{x{\boldsymbol{e}}_{i}}(Z(t)\in A)}{x},\quad t>0.$
Using standard arguments (see, for example, [32, Thm. 8.6]), one can deduce from (2.8) that for $t>0$
@@
Assume that there exists a weak solution to coagulation equation (2.5) with initial condition (2.6). Then it admits the stochastic representation
@@
Remark 2.8.
In the one-dimensional case, it is well known (see, for example, [39, Chap. 4.2, Eq. (4.6)]) that the entrance law of the one-dimensional Feller diffusion is given by
$Q(t)=\frac{2}{t}\,\mathrm{Exp}\!\left(\frac{2}{t}\right),$
where $\mathrm{Exp}(\alpha)$ denotes the exponential distribution with parameter $\alpha>0$. Consequently, in close analogy with the discrete case, one obtains an explicit solution to the continuous coagulation equation (see, for example, [11, 2]) in the form
$u(t,x)=\frac{4}{t^{2}}\exp\Big(-\frac{2x}{t}\Big).$
Moreover, [2, Sec. 3.1] provides an alternative probabilistic interpretation of this result based on a spatial Poisson construction. For each $t>0$, let $\mathcal{P}(t)$ be a Poisson point process on $\mathbb{R}$ with intensity $r(t)\coloneqq 2/t$, coupled by independent thinning so that, for $0<s<t$, $\mathcal{P}(t)$ is obtained from $\mathcal{P}(s)$ by retaining each point with probability $r(t)/r(s)$. This yields a decreasing Markov family of Poisson configurations, defined via an entrance law since $r(t)\to\infty$ as $t\to 0$. The intervals between points form a renewal (Poisson-cluster-type) structure whose lengths represent cluster masses and merge upon point deletion. The induced interval-length distribution then evolves according to the continuous coagulation equation (2.5) for $d=1$.
@@
Assume further that there exists a solution ${{\boldsymbol{u}}}=(u_{i})_{i\in[d]}$ to coagulation equation (2.3) with initial condition (2.4). Then $\boldsymbol{u}$} admits the stochastic representation
@@
It is important to note that, in the critical sampling regime, such a branching representation may fail when the coalescence rates are too small; it holds only for sufficiently large values of $c$, consistent with the large-sampling regime, where $c$ may be regarded as effectively infinite. In the large-sampling regime this phenomenon is particularly striking: regardless of the sampling procedure, the sampled genealogy is a continuous-state branching process whose branching mechanism depends both on the underlying coalescent and on the sampling itself. Although it is well known in the one-dimensional case that coalescents at small times are closely related to branching processes, it was far from clear that this correspondence would extend to higher dimensions.
Let us also observe that the stochastic representations introduced above imply uniqueness of solutions to our coagulation equations under their respective initial conditions. In the critical case, we provide in Appendix A, Lemma A.1, an alternative proof of uniqueness that removes the assumption $d_{i}\geq 0$. These uniqueness results will play a crucial role in establishing the convergence of the empirical measures.
Remark 2.11 (Stochastic representation at equilibrium).
Note that if $\boldsymbol{\beta}=\boldsymbol{\xi}$ is the equilibrium probability measure, i.e.
$\sum_{j\in[d]\setminus\{i\}}\beta_{j}w_{j,i}\;=\;\beta_{i}\sum_{j\in[d]\setminus\{i\}}w_{i,j},\qquad\,i\in[d],$
then, for all $i\in[d]$,
$d_{i}=\frac{c\,\alpha_{i}\beta_{i}}{2}>0,$
and hence the conclusion of the previous theorem holds automatically.
2.6. Conjectures on the site-frequency spectrum
As discussed in the introduction, the original motivation for this work comes from the study of the site-frequency spectrum (SFS) of structured coalescents – a central object in population genetics and a notoriously challenging problem – in the regime of fast migration and large sample sizes. A classical approach in the literature assumes fast migration with a fixed sample size; in this setting, a slow-fast principle implies that the SFS becomes asymptotically indistinguishable from that of a one-dimensional Kingman coalescent, leading to dimension reduction and explicit formulas. While mathematically elegant, this regime is biologically unsatisfactory, since information about the underlying population structure is lost – an effect often referred to as the collapse of structure (see, for instance [14, Chap. 6.3]). Preliminary calculations suggest that the situation changes markedly when the sample size is large and of the same order as, or larger than, the migration scale, which is precisely the regime studied in this paper. We conjecture that in this setting the SFS is directly related to the solution of the coagulation equations introduced above. While a detailed investigation is beyond the scope of the present work, we briefly outline our proposed approach below.
To begin, let $\tau_{K}$ denote the time to the most recent common ancestor (MRCA)
$\tau_{K}\coloneqq\inf\Big\{t\geq 0:\sum_{i\in[d]}L^{K}_{i}(t)=1\Big\}.$
We now define the branch-length measure $B^{K}$ as the random variable valued in ${\mathcal{M}}([N_{K}]^{d}_{0})$ such that
$B^{K}(\{{{\boldsymbol{k}}}\})=\sum_{i\in[d]}\int_{0}^{\tau_{K}}\nu_{i}^{K}(s,\{{{\boldsymbol{k}}}\})\mathrm{d}s,\quad{{\boldsymbol{k}}}\in[N_{K}]^{d}_{0}.$
We also define the rescaled branch-length measure as
${\mathcal{B}}^{K}({\mathrm{d}{\boldsymbol{x}}})=B^{K}({{s_{K}}{\mathrm{d}{\boldsymbol{x}}}})=\sum_{i\in[d]}\int_{0}^{K\tau_{K}}\mu_{i}^{K}(s,{\mathrm{d}{\boldsymbol{x}}})\mathrm{d}s.$
Then Theorem 2.5 already yields
$\sum_{i\in[d]}\int_{0}^{A}\mu_{i}^{K}(s,\mathrm{d}\boldsymbol{x})\,\mathrm{d}s\;\Rightarrow\;\sum_{i\in[d]}\int_{0}^{A}u_{i}(s,\mathrm{d}\boldsymbol{x})\,\mathrm{d}s\quad\text{as }K\to\infty,$
for any fixed time $A>0$. However, to analyze the asymptotic behavior of $\mathcal{B}^{K}$, we must also control the contributions arising after time $A$. We expect these contributions, which are related to the proportion of small blocks, to become negligible as $A$ grows. Making this intuition rigorous, however, would require sharp bounds on the growth of block sizes in the structured coalescent, and establishing such estimates lies beyond the scope of the present work.
In ongoing work, we aim to show the following.
Conjecture 1.
Assume that $\gamma_{K}\to\infty$ as $K\to\infty$. Let $T_{0}=\inf\{t>0:{{\boldsymbol{Z}}}(t)={\bf 0}\}$, where $\boldsymbol{Z}$} denotes the solution of the SDE (2.7), and define $Q_{\beta}=\sum_{i\in[d]}\beta_{i}Q_{i}$. Then
[…]
In other words, $\mathcal{B}$$^{\infty}$ is the potential measure associated to $Q_{\boldsymbol{\beta}}$.
We now assume that neutral mutations (i.e. mutations that do not influence the genealogical dynamics of the coalescent) occur along the branches of the coalescent tree at a constant rate $\theta>0$ (see Figure 3). These mutations are interpreted in the spirit of the infinite-sites model, which assumes that each mutation occurs at a unique site on the genome that has never mutated before. For each $\boldsymbol{k}\in[N_{K}]_{0}^{d}$, a mutation is said to be of type $\boldsymbol{k}$ if it affects $k_{i}$ leaves of color $i$ for all $i\in[d]$. We define the site-frequency measure $S^{K}$ on $N_{K}]_{0}^{d}$ by
$S^{K}(\boldsymbol{k})\coloneqq\#\{\text{mutations of type $\boldsymbol{k}$ occurring before time $\tau_{K}$}\},\qquad\boldsymbol{k}\in[N_{K}]_{0}^{d}.$
Formally, $S^{K}$ is a Poisson point process on $N_{K}]_{0}^{d}$ with intensity measure $\theta\,B^{K}(\mathrm{d}\boldsymbol{x})$.
Figure 3. The coalescent with mutations for $d=3$, $N_{K}=10$. Black circles indicate mutations occurring at constant rate $\theta$ per block.
As a direct consequence of Conjecture 1, we expect the following result to hold.
Conjecture 2.
Assume that $\gamma_{K}\to\infty$ as $K\to\infty$. Then
$S^{K}({\gamma_{K}}{\mathrm{d}\boldsymbol{x}})\Rightarrow\mathrm{PPP}(\theta{\mathcal{B}}^{\infty}(\mathrm{d}\boldsymbol{x}))\quad\text{as }K\to\infty.$
In the case of critical sampling, we anticipate that an analogous result holds.
2.7. Proof strategy for the main results
We begin by recalling that the stochastic representation in Theorem 2.7 was already derived in Section 2.5. Its analogue in the critical regime, stated in Theorem 2.9, will be proved in Appendix A. These results imply that the coagulation equations (2.3) and (2.5), with initial conditions (2.4) and (2.6), respectively, admit at most one (weak) solution; in the critical case, this conclusion requires the additional assumption that $d_{i}>0$ for all $i\in[d]$. In Proposition A.1, we show that uniqueness still holds in the critical case without this additional assumption.
Sections 3, 4, and 5 are devoted to the proofs of Theorems 2.4 and 2.5, which rely on three main components: (1) convergence of the generators, (2) tightness, and (3) characterization of accumulation points.
Part (1) is treated in Section 3, where we also derive moment estimates that are crucial for parts (2) and (3). Section 4 addresses parts (2) and (3). Finally, in Section 5 we complete the proofs of Theorems 2.4 and 2.5. We emphasize that our proof of existence of solutions to the coagulation equations, with their respective initial conditions, is probabilistic: it follows from parts (2) and (3) together with a careful analysis of the asymptotic initial conditions. This analysis is straightforward in the critical regime, but more delicate in the large-sampling regime. The uniqueness results discussed above play a key role in concluding the proofs.
Let us briefly comment on the differences in the analysis between our structured setting and the classical Kingman case. A key insight underlying several arguments in parts (2) and (3) is a coupling that allows us to transfer moment bounds from the classical (non-structured) Kingman coalescent. Beyond this, the structured setting renders parts (1), (2), and (3) technically more involved than in the non-structured case, but without relying on fundamentally new concepts. Nonetheless, the analysis is far from a routine extension: a genuinely new difficulty arises in the study of the initial conditions in the large-sampling regime, where the main challenge is intrinsic to the structured nature of the coalescent.
3. Generator convergence and moment bounds
As announced in the introduction, the proofs of Theorems 2.5 and 2.4 rely on three classical ingredients: (1) convergence of the infinitesimal generators $A^{K}$ of the vector of empirical-measure processes ${\boldsymbol{\mu}}^{K}$, in an appropriate sense; (2) tightness of the sequence of empirical measures; and (3) identification of its accumulation points.
In this section we address (1) (see Proposition 3.2) and establish estimates that will be useful when tackling (2) in the next section. Specifically, we derive $L^{1}$- and $L^{2}$-bounds for the sum of squared block sizes (see Lemma 3.3), and we show that the block-counting process in our model can be sandwiched between the block-counting processes of two Kingman coalescents with different coalescence rates (see Lemma 3.4). This coupling will allow us to carry over moment bounds from the Kingman case, which we state at the end of the section (see Lemma 3.5).